OpenStudy (viixxen):

"A cylinder has a radius of 5x + 2 and a height of 2x + 8 which polynomial in standard form best describes the total volume of the cylinder? Use the formula V=pi r^2h for the volume of a cylinder."

OpenStudy (btaylor):

\[V=\pi r^2 h = \pi (5x+2)^2(2x+8)\]

OpenStudy (btaylor):

Expanding, we get:\[\pi (5x+2)(5x+2)(2x+8) = (50x^3+240x^2+168x+32)\pi\]

OpenStudy (viixxen):

Thank you! (:

OpenStudy (btaylor):

anytime.